1.         What is the change in internal energy of a system that does 7.02 kJ of work and absorbs 888 J of heat?

2.         Given the reaction

                        2 Na2O2(g) + 2 H2O(l) → 4 NaOH(s) + O2(g)            ∆H = -109 kJ

            determine ∆H for the following reactions.

a)16 Na2O2(g) + 16 H2O(l) → 32 NaOH(s) + 8 O2(g)

b)2 NaOH(s) + ½ O2(g) → Na2O2(s) + H2O(l)

please make it in word software
Answer #1

1. ACCORDING TO FIRST LAW OF THERMODYNAMICS

dU = dq + dw = +888- 7020 j =6132 j ( IUPAC Convention: heat absorbtion + ve and work done by system -)

2. a.) delH = 8 x -109 kj = -872 kj

b.) delH = 1/2 x (+109) kj = 54.5 kj